perform sensitivity analysis using PW Method over a range of (+ or -) 20% in
i. initial investment. ii. net annual revenue
iii. salvage value iv. useful life:

Initial investment (Rs) = 200,000
Annual revenue (Rs) =50,000
Annual expenses (Rs) =5,000
Salvage values (Rs) =25,000
Useful life =10 yrs
MARR =12% Per year
Draw also the sensitivity graph.
1
Expert's answer
2020-05-27T10:11:08-0400

"NPV=-200,000+45,000*\frac{1.12^9-1}{1.12-1}+\frac{45,000-25,000}{1.12^{10}}"

"NPV=471,343.99"

i) "I1=240,000\ I2=160,000"

"NPV1=431,343.99\ NPV2=511,343.99"

s-sensitivity

"s1=\frac{(431,343.99-471,343.99)}{471,343.99}*100\%=-9.27\%"

"s2=\frac{511,343.99-471,343.99}{471,343.99}*100\%=9.27\%"

"sensitivity\ NPV=9.27/20*100\%=46.35\%"

1% changes Initial investment gives changes of NPV on 46.35%

ii) "NPV=320,367.7"

"s=-32.03\%"

"sensitivity\ NPV=160.15\%"

iii) "NPV=472,953.86"

"s=0.34%"

"sensitivity\ NPV=1.7%"

iv)"NPV=262,083.19"

"s=-44.4%"

"sensitivity\ NPV=222\%"